OpenStudy (anonymous):

Help! Lim as x approaches pi/6 sinx-1/2/(x-pi/6)

OpenStudy (anonymous):

\[\lim_{x \rightarrow \pi/6} \frac{ sinx-1/2 }{ x-\pi/6 }\]

OpenStudy (anonymous):

use L'Hopitals rule since you have an indeterminate form of 0/0

OpenStudy (psymon):

Doubt she knows l'hopitals.

OpenStudy (anonymous):

Do I start out with sinx-1/2 * sinx+1/2?

OpenStudy (anonymous):

ahhh, basically take derivative of top and derivative of bottom and take that limit

OpenStudy (psymon):

Yeah, but assuming shes not supposed to know l'hopitals, we dont want to just teach it, shes supposed to find it a different way.

OpenStudy (anonymous):

using l hospitals rule, cosx-0/1 now put the limit value in place of x. cos pie/6 root3/2

OpenStudy (anonymous):

cos x/1 then it's just cos pi/6 which is sqrt3/2

OpenStudy (anonymous):

I dont think you can take the limit without it though

OpenStudy (anonymous):

we nicknamed lhopitals the limit breaker lol

OpenStudy (anonymous):

Oh, we didn't learn that yet.

OpenStudy (anonymous):

this might be the problem to get you there

OpenStudy (psymon):

I have a suggestion then, thats not l'hopitals. Doesnt mean itll work, but....

OpenStudy (anonymous):

u havent learned l hospitals?then first go through the theory part in ur text book.

OpenStudy (anonymous):

its pretty easy, if you get a form of 0/0 or infinite/infinite you can use lhopitals rule

OpenStudy (anonymous):

if you take the limit of that then that's the limit for the function

OpenStudy (anonymous):

Do I just times the bottom and top by sinx+1/2? Will that work?

OpenStudy (anonymous):

that might you can always multiply a function by 1 in any form

OpenStudy (anonymous):

but it really just makes life hard lol

OpenStudy (anonymous):

Well, that's the way the professor wants it lol.

OpenStudy (anonymous):

in l hospitals rule,first u have to check weather the numerator and denominator both are coming to zero by putting the value of limit in place of x.

OpenStudy (anonymous):

Of course the denominator has a zero.

OpenStudy (anonymous):

yea if that's the case then manipulate the equation to a form that you can plug in the limit for without dividing by 0,

OpenStudy (anonymous):

which class?

OpenStudy (anonymous):

Calculus I

OpenStudy (anonymous):

ahhh yea you learn L'hopitals in Calc 2

OpenStudy (anonymous):

you need it to break open harder limits

OpenStudy (anonymous):

The other way is to put the function in your calculator and get ever so close to pi/6 but not at pi/6

OpenStudy (anonymous):

We have to prove it though by writing it out.

OpenStudy (anonymous):

basically however you do it the answer is sqrt3/2 so if you get that you're good. =)

OpenStudy (anonymous):

Actually, my question is how do you get to that answer.. but thanks lol

OpenStudy (psymon):

No, it wont. You HAVE to get rid of the x-pi/6 deal, that won't go away. Like I said, a suggestion that may or may not work. So, substitution time. \[u = x-\frac{ \pi }{ 6 }\] therefore \[x = u+ \frac{ \pi }{ 6 }\] \[\frac{ \sin(u+\frac{ \pi }{ 6 })-\frac{ 1 }{ 2 } }{ u }\] Use sum of sines formula to expand the sin portion: \[\frac{ sinucos(\frac{ \pi }{ 6 })+\sin(\frac{ \pi }{ 6 })cosu-\frac{ 1 }{ 2 } }{ u }\] \[\frac{ \frac{ \sqrt{3} }{ 2 }sinu+\frac{ 1 }{ 2 }cosu-\frac{ 1 }{ 2 } }{ u }\] Now because of theu-substitution, as x goes to pi/6, u goes to 0. So we have to keep this in mind as I finish. Now I'll split this into two fractions: \[\lim_{u \rightarrow 0}\frac{ \frac{ \sqrt{3} }{ 2 }sinu }{ u }+\frac{ \frac{ 1 }{ 2 }(cosu-1) }{ u }\]

OpenStudy (psymon):

Now there are two limit identities that can be used when the limit approaches 0. The first is \[\lim_{x \rightarrow 0}\frac{ sinx }{ x }= 1\] the second is \[\lim_{x \rightarrow 0}\frac{ cosx -1 }{ x }= 0\]Well, looks like to me we have both of those identities in our two fractions. So as u goes to o, that left fraction that has sinu/u goes to 1. And the right fraction that has (cosu-1)/u becomes 0. SO that means you're left finally with \[\lim_{u \rightarrow 0}\frac{ \sqrt{3} }{ 2 }(1) + \frac{ 1 }{ 2 }(0) = \frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

Thank you so much! I actually understand it now. I forgot that I had to make u=x-a and x=u+a as x approaches a and u approaches 0. You get best response!

OpenStudy (psymon):

Its a really sneaky substitution. At first glance, the calc 2 method just looks like it may be the only way, but you can sometimes manipulate things in some awesome ways. Glad you understand what I did ^_^

OpenStudy (anonymous):

Could you help me with one more thing by the way? How would I explain my answer for sin (.53)-1/2 over .53 - pi/6 is .86442 without using a calculator?

OpenStudy (anonymous):

They gave me a hint that said pi/6 = .523598

OpenStudy (psymon):

It looks like the exact same problem.

OpenStudy (anonymous):

Well nevermind, you helped enough lol. It says multiple choice (do not use calculator) then put the problem up and it says explain your answer. I don't know what's with that.

OpenStudy (psymon):

Well, you just would rewrite it into the same problem you had before. You just would let .53 be x and say its a limit as x goes to .53, then do the same thing.

OpenStudy (anonymous):

Oh okay! That makes sense lol thank you.

OpenStudy (psymon):

yeah, sure xD